# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ Abstract
We revisit the theory of Kac–Moody symmetric spaces in order to include the affine case, complex numbers, and Galois descent to almost split real Kac–Moody symmetric spaces.
@@ 1 Introduction
Kac–Moody groups have been studied extensively since the 1980s, initiated by [KP85] and [Tit87]. However, for some reason Kac–Moody symmetric spaces have only been formally introduced in [FHHK20], although the heuristic idea that Kac–Moody symmetric spaces might be useful has been around in particular in mathematical physics for some time.
The focus of [FHHK20] was on symmetric spaces for split real Kac–Moody groups modelled with maximal flats of minimal possible dimension in order to propose geometries that might allow a primitive action as in [Cap05] – as this might then help prove abstract simplicity of the acting group. It is thus not surprising that the methods of [FHHK20] failed for affine Kac–Moody groups.
In the present article we vary on the theme of Kac–Moody symmetric spaces and introduce models with higher-dimensional flats that finally provide examples of Kac–Moody symmetric spaces also in the affine case. The key difference in the modelling can actually be found already on Kac–Moody algebra level and is described in Remark 3.3 and Convention 3.4. Not surprising, it hence turns out that the Kac–Moody symmetric spaces introduced here are identical to the ones introduced in [FHHK20] if and only if the generalized Cartan matrix is invertible (see Section 7).
Moreover, in the present article we also include the field $\mathbb{C}$ of complex numbers (see Convention 3.5), and we conduct Galois descent based on [Rém02], in order to model Kac–Moody symmetric spaces also in the almost split real case; see Section 9.
We took quite some effort to make the present article accessible independent from [FHHK20], but it is quite natural that we need to borrow and cite a wealth of results from [FHHK20]; on top of that, also a detailed comparison of our results with the results from [FHHK20] is called for.
We stress again that the purpose of this article is to provide an alternative definition in order to approach Kac–Moody symmetric spaces, that allows to also cover the affine case. On top of that we deal with the complex case and the almost split real case.
Our main result is the following:
Theorem 9.2
Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and let $G$ be an (almost) split Kac–Moody $\mathbb{K}$-group satisfying (8.2) and denote by $G(\mathbb{K})$ the rational points of $G$. The quotient
@@
is a symmetric space in the sense of Loos [Loo69] with the reflection map $\mu^{\operatorname{as}}$. Moreover, there is a natural action of $G(\mathbb{K})$ on the symmetric space given by automorphisms
@@ Acknowledgment:
The authors thank the DFG SPP 2026 priority program “Geometry at infinity” for partial financial support via KO 4323/14. The first half of the project, both authors were affiliated with and funded by JLU Giessen, the second half affiliated with and funded by CAU Kiel. The authors would also like to thank Paul Zellhofer in particular for a wealth of very helpful discussions and comments. This article is based on the PhD thesis of the second author that has been advised by the first author.
@@ 2 Basics on Symmetric Spaces
    for every $x\in X$ there is a neighborhood $U$ such that $x.y=y$ implies $x=y$ for all $y\in U$.
@@
$\begin{matrix}\textbf{S.4}_{\text{global}}&\forall x,y\in X\;\;y=y\Longrightarrow x=y.\end{matrix}$
@@
Remark 2.4
Consider a pointed symmetric space $(X,\mu)$ with base point $b\in X$. We now want to describe the symmetric space via the quadratic representation, see [Loo69, Section II.1] or [FHHK20, Remark 2.10]. For $x\in X$ define
@@
is surjective. Since $t_{x}\circ s_{b}=s_{x}$ and since $s_{x}=s_{y}$ implies $x=y$, the map $\eta$ is also injective.
Model a symmetric space on the set $T(X,b)$ by defining the reflection map
$\mu_{T}\colon T(X,b)\times T(X,b)\to T(X,b)\,,\;(s,t)\mapsto s\circ t^{-1}\circ s.$
A direct calculation shows that $\eta:(X,\mu)\to(T(X,b),\mu_{T})$ is an isomorphism of symmetric spaces. Therefore, one has to check that the reflection map of $(X,\mu)$ is preserved, i.e. $\eta(\mu(x,y))=\mu_{T}(t_{x},t_{y})$. On the right-hand side one can calculate
@@
Recall from [FHHK20, Section 2.15] the following useful definitions and properties of an abstract symmetric space $(X,\mu)$:
@@
  • Call a subspace $U\subseteq X$
@@
  • Two points $x,y\in X$ weakly commute, if for every point $p\in X$ applies $x.(p.(y.p))=y.(p.(x.p))$.
  • Two points $x,y\in X$ commute, if for all points $p,q\in X$ applies $x.(p.(y.q))=y.(p.(x.q))$.
@@
Definition 2.5 (cp. Definition 2.23 from [FHHK20])
@@ 3.1 Kac–Moody Algebra
According to the approach proposed in Kac’s book [Kac90, §1.1], one wants the sets $\Pi$ and $\Pi^{\vee}$ to be linearly independent. Therefore, in the case that $\mathbb{A}$ is of affine type it follows from the relation $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=a_{ij}\in\mathbb{A}$, that $\Pi$ is linearly dependent. Hence, extending the Cartan subalgebra $\mathfrak{h}$ solves this problem, as done in [Kac90, Proof of Proposition 1.1] and [Mar18, Chapter 3.5 and Example 7.10]: Let $\mathbb{A}\in\mathbb{Z}^{n\times n}$ and $l$ be the rank of $\mathbb{A}$, then the extended Cartan subalgebra must be $\mathfrak{h}\cong\mathbb{C}^{2n-l}$. The dual space is also extended and constructed so that the roots still satisfy the relation to $\mathbb{A}$.
Let us now denote by $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ the root datum where the elements $c_{i}$ and $h_{i}$ are linearly independent and denote by $\Lambda_{\mathrm{Kac}}$ the corresponding $\mathbb{Z}$-module, which has minimum rank with respect to these properties, i.e. the rank of $\Lambda_{\mathrm{Kac}}$ is equal to $2n-l$. Similarly, we use $\Lambda_{\mathrm{Kac}}^{\vee}$ to denote the $\mathbb{Z}$-dual. Moreover, denote by $\{v_{i}\mid 1\leq i\leq 2n-l\}$ the $\mathbb{Z}$-basis of the extended $\Lambda_{\mathrm{Kac}}$, where $h_{i}=v_{i}$ for $1\leq i\leq n$. We follow here the notation and idea of [Mar18, Example 7.10].
@@
for $h\in\mathfrak{h}_{\mathcal{D}}$ and $i,j\in I$. In the following, we denote with $\mathfrak{g}(\mathbb{A})$ the Kac–Moody algebra of type $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$.
@@
For the rest of the article we consider the root datum $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$, where $\mathbb{A}$ is a indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$.
@@
Denote by $\Delta$ all roots of $\mathfrak{g}(\mathbb{A})$. Furthermore, a root $\alpha$ is called positive if the coefficients in $Q$ are non-negative, and negative otherwise. By [Mar18, Proposition 3.14 (5)] follows that every root is either positive or negative.
@@ 3.2 The Tits Functor
Let $R$ be a ring and recall that $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}=\left(I,\mathbb{A},\Lambda_{\mathrm{Kac}},(c_{i})_{i\in I},(h_{i})_{i\in I}\right)$. Then a split torus scheme (cp. [Mar18, Section 7.4]) is given by
$T\colon\mathbb{Z}\text{-alg}\to\textbf{Grp}\,,\;T(R)\coloneqq\Lambda_{\mathrm{Kac}}^{\vee}\otimes_{\mathbb{Z}}R^{\times},$
@@
    For every ring $R$, the homomorphism $\eta\colon\mathfrak{T}_{\Lambda}(R)\to\mathcal{G}(R)$ is injective.
@@
Note, that on fields, the Tits functor is unique up to isomorphism, see [Mar18, Theorem 7.82].
@@
For the remainder of this article, let $\mathbb{K}\in{\mathbb{R},\mathbb{C}}$ and denote by $G$ the evaluation of the Tits functor $\mathcal{G}$ with respect to the Kac–Moody root datum $\mathcal{D}_{\text{Kac}}^{\mathbb{A}}$ on the field $\mathbb{K}$; $G$ is called a split (minimal) Kac–Moody group over $\mathbb{K}$ (of type $\mathbb{A}$) and $T$ its (standard) torus of $G$.
@@
Further define for $\alpha\in\Delta^{\mathrm{re}}$ the corresponding rank one subgroup of $G$ as $G_{\alpha}\coloneqq\langle U_{\alpha},U_{-\alpha}\rangle$. Note that $G_{\alpha}\cong\operatorname{SL}_{2}(\mathbb{K})$, see [FHHK20, Proposition 3.17].
Define the (standard) Borel subgroups of $G$ by $B_{\pm}\coloneqq T\ltimes U_{\pm}$ (cp. [AB08, Corollary 7.122]) and denote all their conjugates as Borel subgroups.
@@
  • $G$ is a $k_{\omega}$-group ([HKM13, Proposition 7.10] and [HK23, Definition and Remark 2.1]).
  • The Kac–Peterson topology induces the unique Lie group topology on $T$, in particular $T$ is a closed subgroup ([HK23, Definition and Remark 2.1] and [HKM13, Corollary 7.17]).
@@
Now we prove theorem [FHHK20, Proposition 3.10] in the context of this article. This allows us to provide the torus with a Lie group topology.
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ and rank $l$. Further denote by $\mathfrak{g}(\mathbb{A})$ the corresponding Kac–Moody algebra. Equip $\mathfrak{h}\cong\mathbb{K}^{2n-l}$ with the vector space topology and $T\cong(\mathbb{K}^{\times})^{2n-l}$ with the Kac–Peterson topology. Then the exponential map
$\exp_{\operatorname{KM}}\colon\mathfrak{h}\to T$
is a quotient map. In particular $T$ is a Lie group homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ equipped with its standard topology.
@@
where $J=\{1,\ldots,2n-l\}$ and by construction follows that $\langle u_{j},v_{i}\rangle=b_{ij}$ and that $\langle u_{j},v_{i}\rangle=a_{ji}$ for $i,j\in I$. Denote by $\widetilde{G}$ the Kac–Moody group associated to $\widetilde{\mathcal{D}}$.
@@
where $\alpha_{i}=c_{i}\otimes 1$ for $i\in I$. Compare this to [HK23, Proposition B.5] or [Voc23, Appendix C], where it is done for arbitrary split Kac–Moody groups.
Both are Kac–Moody groups and since we assumed that the Tits functor is non-degenerate, it is already clear that the relations of a Kac–Moody group (see [Mar18, Definition 7.47]) are preserved under the mapping $\pi$. Moreover, both groups can be equipped with the Kac–Peterson topology, and by construction is $G$ a subgroup of $\widetilde{G}$. Based on the construction, $\mathbb{B}$ is invertible and thus $\widetilde{\mathcal{D}}$ has finite center. This allows us to apply [HKM13, Proposition 7.18]. It follows from this proposition that the torus of $\widetilde{G}$ endowed with the Kac–Peterson topology is homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ equipped with its standard topology. Moreover the torus of $\widetilde{G}$ equals the torus of $G$ by construction of the root data. ∎

### v4

# Kac–Moody Symmetric Spaces: arbitrary symmetrizable complex or almost split real type

@@ Abstract
We revisit the theory of Kac–Moody symmetric spaces (of ”non-compact type”) in order to include the affine case, complex numbers, and Galois descent to almost split real Kac–Moody symmetric spaces.
@@ 1 Introduction
Non-compact (semi)simple Lie groups admit two natural geometries: a symmetric space of non-compact type and a Tits building in the boundary at infinity of the symmetric space; cf. [BS87]. From a purely group-theoretic point of view, the symmetric space is suitable for establishing, resp. confirming the simplicity of the acting Lie group as an abstract group – the simplicity is related to the maximality of the maximal compact subgroup (that is, the maximal compact subgroup actually being a maximal subgroup) with in turn is related to the primitivity of the Lie group action on the symmetric space; cf. [Cap05]. Kac–Moody groups have been extensively studied since the 1980s, initiated by [KP85] and [Tit87], and it turns out that they actually admit two buildings which form a so-called twin building. However, Kac–Moody symmetric spaces have only been formally defined and investigated in 2020 in [FHHK20] and only over the real numbers and only for indefinite type, although the heuristic idea that Kac–Moody symmetric spaces might be useful has been around in particular in mathematical physics for some time; cf. [DHN03, Section 8.3]. Kac–Moody symmetric spaces actually admit a causal structure defining a causal partial order ([FHHK20, Proposition 7.46] combined with [Zel25, Theorem 6.3.18]). The twin building can be recovered from inside the causal boundary of the Kac–Moody symmetric space; cf. [FHHK20, Corollary 7.41].
It is still an open question whether split real Kac–Moody groups of indefinite type are abstractly simple. One of the goals of [FHHK20] was to propose a symmetric space that might allow a primitive action as in [Cap05] – in order to establish the abstract simplicity of the acting group. In case of non-invertible generalized Cartan matrix this lead to some modeling choices that eventually excluded affine types – which is not surprising as affine Kac–Moody groups are well known to not be abstractly simple.
In the present article we make different modeling choices. For split real Kac–Moody groups with invertible generalized Cartan matrix they lead to exactly the same class of symmetric spaces – see also the discussion in Section 7 below. For a non-invertible generalized Cartan matrix these modeling choices lead to a symmetric space which obviously is imprimitive, and they allow to also consider the case of an affine generalized Cartan matrix. Moreover, in the present article we also include the field $\mathbb{C}$ of complex numbers (see Convention 3.5), and we conduct Galois descent based on [Rém02], in order to model Kac–Moody symmetric spaces also in the almost split real case; see Section 9.
The following result is our main existence theorem.
Proposition 4.4 and Theorem 9.2
Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and let $G$ be a split or almost split Kac–Moody $\mathbb{K}$-group (satisfying (8.2) in the almost split situation) and denote by $G(\mathbb{K})$ the $\mathbb{K}$-rational points of $G$. Let $\Theta$ be the Cartan–Chevalley involution of $G(\mathbb{K})$ over the real numbers, resp. the Cartan–Chevalley involution composed with the continuous field involution over the complex numbers, and let $K$ be the subgroup of $G$ of $\Theta$-fixed elements. Then the quotient
@@
is a symmetric space in the sense of Loos [Loo69] with respect to the reflection map
$(gK(\mathbb{K}),hK(\mathbb{K}))\mapsto g\Theta(g)^{-1}\Theta(h)K(\mathbb{K}).$
Moreover, there is a natural action of $G(\mathbb{K})$ on the symmetric space given by automorphisms
@@
The condition (8.2) is only relevant for the almost split situation and is central to the theory of almost split Kac–Moody groups: it guarantees that the rational points of the Kac–Moody group correspond to the fixed points of the complex Kac–Moody group under Galois involution. For details on this involution see section 8.1 and for details on the condition see section 8.2.
In analogy to [FHHK20] we also study flats and automorphisms of these symmetric spaces and the Weyl group action. We took quite some effort to make the present article accessible independently from [FHHK20], but it is quite natural that we need to borrow and cite a wealth of results from [FHHK20]; on top of that, also a detailed comparison of our results with the results from [FHHK20] is called for.
@@ Acknowledgment:
The authors thank the DFG SPP 2026 priority program “Geometry at infinity” for partial financial support via KO 4323/14. During the first half of the project both authors were affiliated with and funded by JLU Gießen, during the second half affiliated with and funded by CAU Kiel. The authors would also like to thank Paul Zellhofer in particular for a wealth of very helpful discussions and comments. Moreover, the authors thank an anonymous referee for literally hundreds of valuable observations and comments that tremendously helped improve the article. This article is based on the PhD thesis of the second author that has been advised by the first author.
@@ 2 Basics on Symmetric Spaces
    for every $x\in X$ there is a neighborhood $U$ of $x$ such that $x.y=y$ implies $x=y$ for all $y\in U$.
@@
$\begin{matrix}\textbf{S.4}_{\text{global}}&\forall x,y\in X\;\;x.y=y\Longrightarrow x=y.\end{matrix}$
@@
Consider a pointed symmetric space $(X,\mu)$ with base point $b\in X$. We now want to describe the symmetric space via the quadratic representation, see [Loo69, Section II.1] or [FHHK20, Remark 2.10].
For $x\in X$ define
@@
By defining the following map
$\mu_{T}\colon T(X,b)\times T(X,b)\to T(X,b)\,,\;(s,t)\mapsto s\circ t^{-1}\circ s,$
the set $T(X,b)$ is turned into a symmetric space with reflection map $\mu_{T}$.
Proposition 2.4
The symmetric spaces $\left(T(X,b),\mu_{T}\right)$ and $(X,\mu)$ are isomorphic.
Proof:
@@
[…]
@@
Recall the following concepts of an abstract symmetric space $(X,\mu)$ from [FHHK20, Section 2.15]:
@@
  • Call a subset $U\subseteq X$
@@
  • Two points $x,y\in X$ weakly commute, if for every point $p\in X$ we have: $x.(p.(y.p))=y.(p.(x.p))$.
  • Two points $x,y\in X$ commute, if for all points $p,q\in X$ we have: $x.(p.(y.q))=y.(p.(x.q))$.
@@
Definition 2.5 (cf. Definition 2.23 from [FHHK20])
@@ 3.1 Kac–Moody Algebra
The original approach of Kac is to define that the sets $\Pi$ and $\Pi^{\vee}$ are linearly independent, see [Kac90, §1.1]. In the general case where the generalized Cartan matrix $\mathbb{A}$ is not necessarily invertible (e.g. if $\mathbb{A}$ is of affine type), we have that $\Pi$ is not linearly independent. To address this problem, one can extend the Cartan subalgebra $\mathfrak{h}$ to ensure that the $\alpha_{i}$ are linearly independent. For a concrete discussion of this, see [Kac90, Proof of Proposition 1.1] and [Mar18, Chapter 3.5 and Example 7.10]. It turns out that for $\mathbb{A}\in\mathbb{Z}^{n\times n}$, where $l$ denotes the rank of $\mathbb{A}$, the extended Cartan subalgebra must have dimension $2n-l$ and is isomorphic to $\mathbb{C}^{2n-l}$. The dual space is also extended and constructed in such a way that the simple roots still have the relation $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=a_{ij}\in\mathbb{A}$.
On this basis, we denote the following as $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ the root datum where the elements $c_{i}$ and $h_{i}$ are linearly independent and denote by $\Lambda_{\mathrm{Kac}}$ the corresponding $\mathbb{Z}$-module, which has minimum rank with respect to these properties, i.e. the rank of $\Lambda_{\mathrm{Kac}}$ is equal to $2n-l$. Similarly, we use $\Lambda_{\mathrm{Kac}}^{\vee}$ to denote the $\mathbb{Z}$-dual. Moreover, denote by $\{v_{i}\mid 1\leq i\leq 2n-l\}$ the $\mathbb{Z}$-basis of the extended $\Lambda_{\mathrm{Kac}}$, where $h_{i}=v_{i}$ for $1\leq i\leq n$. We follow here the notation and idea of [Mar18, Example 7.10].
@@
for $h\in\mathfrak{h}_{\mathcal{D}}$ and $i,j\in I$. Note that $\delta^{i}_{j}$ here means the Kronecker symbol. In the following, we denote with $\mathfrak{g}(\mathbb{A})$ the Kac–Moody algebra of type $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$.
@@
For the remainder of the article we consider the root datum $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$.
@@
In [Mar18, Chapter 3.4] the concept of gradation is explained, which describes the representation of a vector space as the direct sum of subspaces. Following this, this root lattice is called a $Q$-gradation of the root space.
Denote by $\Delta$ all roots of $\mathfrak{g}(\mathbb{A})$. Furthermore, a root $\alpha$ is called positive if the coefficients this representation are non-negative, and negative otherwise. By [Mar18, Proposition 3.14 (5)] follows that every root is either positive or negative.
@@ 3.2 The Tits Functor
Let $R$ be a ring and recall that $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}=\left(I,\mathbb{A},\Lambda_{\mathrm{Kac}},(c_{i})_{i\in I},(h_{i})_{i\in I}\right)$. Then a split torus scheme (cf. [Mar18, Section 7.4]) is given by
$T\colon\mathbb{Z}\text{-alg}\to\textbf{Grp}\,,\;T(R)\coloneqq\Lambda_{\mathrm{Kac}}^{\vee}\otimes_{\mathbb{Z}}R^{\times};$
@@
    For every ring $R$, the homomorphism $\eta\colon T(R)\to\mathcal{G}(R)$ is injective.
@@
Note that, on fields, the Tits functor is unique up to isomorphism, see [Mar18, Theorem 7.82].
@@
For the remainder of this article, let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and denote by $G$ the evaluation of the Tits functor $\mathcal{G}$ with respect to the Kac–Moody root datum $\mathcal{D}_{\text{Kac}}^{\mathbb{A}}$ on the field $\mathbb{K}$; $G$ is called a split (minimal) Kac–Moody group over $\mathbb{K}$ (of type $\mathbb{A}$) and $T$ its (standard) torus of $G$.
@@
Further define for $\alpha\in\Delta^{\mathrm{re}}$ the corresponding rank one subgroup of $G$ as $G_{\alpha}\coloneqq\langle U_{\alpha},U_{-\alpha}\rangle$. Note that by the axiom KMG.3 the matrix
$\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}$
is mapped by $\varphi_{i}$ to $(-1)^{h_{i}}$, which is not the trivial element of the corresponding rank one subgroup, see [Mar18, Exercise 7.33], [Rém02, Chapter 8.2.1] and [Rém02, Chapter 8.4.2]. Therefore $G_{\alpha}\cong\operatorname{SL}_{2}(\mathbb{K})$, for details see [FHHK20, Proposition 3.17].
Define the standard Borel subgroups of $G$ by $B_{\pm}\coloneqq T\ltimes U_{\pm}$ (cf. [AB08, Corollary 7.122]) and the conjugates of these groups are called Borel subgroups.
@@
  • The Kac–Peterson topology induces the unique Lie group topology on $T$ ([HK23, Definition and Remark 2.1] and [HKM13, Corollary 7.17]).
@@
Now we prove [FHHK20, Proposition 3.10] in the context of this article. This allows us to endow the torus with a Lie group topology.
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ and rank $l$. Further denote by $\mathfrak{g}(\mathbb{A})$ the corresponding Kac–Moody algebra. Equip $\mathfrak{h}\cong\mathbb{K}^{2n-l}$ with the vector space topology and $T\cong(\mathbb{K}^{\times})^{2n-l}$ with the Kac–Peterson subspace topology. Then $T$ is a Lie group homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ equipped with its standard topology.
@@
where $J=\{1,\ldots,2n-l\}$ and by construction we have $\langle u_{j},v_{i}\rangle=b_{ij}$ and that $\langle u_{j},v_{i}\rangle=a_{ji}$ for $i,j\in I$. Denote by $\widetilde{G}$ the Kac–Moody group associated to $\widetilde{\mathcal{D}}$.
@@
where $\alpha_{i}=c_{i}\otimes 1$ for $i\in I$. Compare this to [HK23, Proposition B.5] or [Voc23, Appendix C], where it is done for arbitrary split Kac–Moody groups. As in [Har20, Proposition 3.5.2], the kernel of $\pi$ is a finite central subgroup, possibly trivial.
Both are Kac–Moody groups and since we assumed that the Tits functor is non-degenerate, it is already clear that the relations of a Kac–Moody group (see [Mar18, Definition 7.47]) are preserved under the mapping $\pi$. Moreover, both groups can be equipped with the Kac–Peterson topology, and by construction is $G$ a subgroup of $\widetilde{G}$. Based on the construction, $\mathbb{B}$ is invertible and thus $\widetilde{\mathcal{D}}$ has finite center (see [Har20, Proposition 3.5.2]). This allows us to apply [HKM13, Proposition 7.18]. It follows from this proposition that the torus of $\widetilde{G}$ endowed with the Kac–Peterson topology is homeomorphic to $(\mathbb{K}^{\times})^{2n-l}/\ker(\pi)$ equipped with its standard Lie topology. Note, that $\ker(\pi)$ is a finite central subgroup by [Har20, Proposition 3.5.2]. We conclude that the torus of $G$ is homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ as a Lie group, covering the Lie group $(\mathbb{K}^{\times})^{2n-l}/\ker(\pi)$. ∎